An airliner must reach a speed of 110m/s to take off. If the available runway is 2.4 km and the aircraft accelerates uniformly from rest at one end, what minimum acceleration must be available if it is to take off?

1
Expert's answer
2022-02-11T08:48:31-0500

Given:

"v_f=110\:\rm m/s"

"v_i=0\:\rm m/s"

"d=2400\:\rm m"


The distance traveled by aircraft within runway 

"d=\frac{v_f^2-v_i^2}{2a}"

Hence, the acceleration

"a=\frac{v_f^2-v_i^2}{2d}=\frac{110^2-0^2}{2*2400}=2.52\:\rm m/s^2"